The wavelength of light used in Young’s double slit experiment is 6000 Å. The 

second and the fourth bright fringes from the centre of the fringe pattern are 

located respectively at 10.24 mm and 12.40 mm. If the observation screen is placed 

at a distance of 1 m from the slits, calculate the separation between the slits.


1
Expert's answer
2021-06-07T09:29:23-0400

"x_2=\frac{2\lambda l}{d},"

"x_4=\frac{4\lambda l}{d},"

"d=\frac{n\lambda l}{x_n},"

"d=\frac{2\cdot 4\lambda l}{x_2\cdot x_4}=37.8~mm."